OpenStudy (anonymous):

Variable a is 4 more than variable b. Variable a is also 1 less than b. Which of the following pairs of equations best models the relationship between a and b? a = b + 4 a = b − 1 a = 4b a = b + 1 b = a + 4 b = a − 1 b = 4a b = a + 1

OpenStudy (anonymous):

1 . a = b + 4 a = b − 1 2. a = 4b a = b + 1 3. b = a + 4 b = a − 1 4. b = 4a b = a + 1

OpenStudy (anonymous):

@iGreen

Nnesha (nnesha):

is means equal sign more than + sign less than - sign( switch variables for example x less than y = y -x)

Nnesha (nnesha):

 write an equation